If -32 is subtracted from a number the result is -52. Find the number
Answer by ilivea9(5)

-32 subtracted from a number (define as "X"), equals -52
Solve for X to find the number.

X - (-32) = -52
X = -52 + (-32)
X = -84

or

X - (-32) = -52
X + 32 = -52
X = -52 - 32
X = -84